ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 90 Yunusi, Mahmadyusuf, Tajik State National University, Tajikistan Mathematical Models of workers potential function and some its applications The report is devoted to construction and investigation of workers potential function. It is showed that this function is a solution of special class differential equation and it is represented in the form of

φ(a) + \frac{∫_{a}^{a_{max}} B(ξ) e^{δ(ξ−a)}dξ}{∫_{a_{min}}^{a_{max}} ∫_{a}^{a_{max}} B(ξ) e^{δ(ξa)}dξda}

where a is an age of workers, B(a) is the fatigue function. δ is a maximal real root of the equation

∫_{0}^{a max} B(ξ)e^{−δξ}dξ = 1 .

Such approach explains many unknown characteristics of economical processes (for example, oscillating character of workers number, an age factor in production processes and others).